OpenStudy (anonymous):

does anyone have the answer? Evaluate 4a+2b-3c if a=2, b=3, and c=4

OpenStudy (saifoo.khan):

2

OpenStudy (bobbyleary):

4(2)+2(3)-3(4) = 8 + 6 - 12 = 2

OpenStudy (anonymous):

put the numbers in place of a, b and c in the polynomial, carry out the multiplication first and then solve it...